# FlexVAR: Flexible Visual Autoregressive Modeling without Residual Prediction

### v1

@@ 4.2 State-of-the-art image generation
BigGAN [3] $\times$ 8.43 177.9 112M
ADM [10] $\times$ 23.24 101.0 554M
DiT-XL/2 [32] $\times$ 3.04 240.8 675M
MaskGIT [5] $\times$ 7.32 156.0 227M
VQGAN [12] $\times$ 26.52 66.8 1.4B
VAR-$d36$ [12] $\times$ 2.63 303.2 2.3B
@@
Table 3: Zero-shot inference on ImageNet 512$\times$512 conditional generation. Training Free indicates whether the model is trained at the 512$\times$512 resolution.
Zero-shot inference on ImageNet 512$\times$512 benchmark. We use FlexVAR-$d24$ to generate 512$\times$512 images and evaluate on ImageNet-512 benchmark without training, as shown in Tab. 3. Surprisingly, our FlexVAR-$d24$ exhibits competitive performance when compared to VAR, despite FlexVAR being trained only on resolutions $\leq$ 256$\times$256 and having only 1.0B parameters.
@@ 4.3 Ablation study
We conduct ablation studies on various design choices in FlexVAR . Due to the limited computational resources, we report the results trained with a short training scheme in Tab. 4, 5, 6, i.e., 40 epochs ($\sim$ 70K iterations).
Component-wise ablations. To understand the effect of each component, we start with standard VAR and progressively add each design. (Tab. 4):
  • Baseline: VAR uses a residual prediction paradigm and exhibits decent performance (1st result), but its flexibility in image generation does not meet expectations (as described in Sec. 1).
  • Position embedding: The introduction of our scalable PE (last result) provides high flexibility for image generation, and further enhances the performance to 3.71 FID.
Figure 3: Compared with VQVAE tokenizers [44, 41] for multi-scale reconstructing images, we downsample the latent features in VQVAE to multiple scales and then use the VQVAE Decoder to reconstruct images. We upsample images $<$ 100 pixels using bilinear interpolation for a better view.
Reconstruct images with different VQVAEs. In Fig. 3, we reconstruct multi-scale images by scaling the latent features in VQVAE tokenizers. Existing VQVAE tokenizers typically do not support scaling the latent features across a range of small to large scales. VAR’s VQVAE [44] uses a residual-based training recipe, directly applying it to non-residual image reconstruction does not yield the anticipated results (the 1st row). The VQVAE tokenizer from Llamagen [41] shows excellent reconstruction performance only at the original latent space, indicating that it is not feasible for scale-wise autoregressive modeling (the 2nd row).
Transfer FlexVAR to Mamba. Recent work, AiM [23], uses the Mamba architecture for token-wise autoregressive modeling. Inspired by this, we modify FlexVAR with Mamba to evaluate the performance (Tab. 5). With similar model parameters, Mamba demonstrates competitive results compared to transformer models, indicating the GT prediction paradigm can effectively adapt to linear attention mechanisms like Mamba. However, considering that this Mamba architecture does not reflect the speed advantage, we do not integrate Mamba into our final version.
Mamba’s inherent unidirectional attention mechanism prevents image tokens from achieving global attention within the same scale. To address this issue, we employ 8 scanning paths in different Mamba layers to capture global information. The specific Mamba architecture is detailed in the Appendix.
@@
Table 5: Ablation of the Mamba architectural.
@@
Table 6: Ablation of Position Embedding. We report the results with model scale -$d20$ trained 40 epochs ($\sim$ 70K iterations). $\times$ denotes that the corresponding Position Embedding is removed.
Position Embedding. In Tab. 6, we experiment with several types of step PE and x-y coordinate PE. To make the model robust to inference steps and enable it to generate images at any resolution, we remove the fixed-length step embedding (results in the second row), and the performance showed only slight changes. We adopt a non-parametric variant, similar to ViT [11], which shows a 0.03 FID difference compared to the learnable variant.
4.4 Analysis of the GT prediction paradigm
Convergence rate. We compared the training loss of VAR and our FlexVAR, as shown in Fig. 4. FlexVAR demonstrates significantly lower loss values and faster convergence, indicating that predicting ground-truth rather than residuals is more friendly for training. This may be attributed to the residuals at different scales lacking semantic continuity, and this implicit prediction approach might limit the training convergence rate.
Figure 4: Training loss of VAR vs. FlexVAR. FlexVAR demonstrates a faster convergence rate. We report the results with trained 40 epochs ($\sim$ 70K iterations).
Generate images at any resolution. We show generated images at different resolutions using FlexVAR-$d24$ in Fig 1 and Fig. 5. By controlling the inference steps, our FlexVAR can generate images at any resolution, despite being trained only on images with resolutions $\leq$ 256px. The generated images demonstrate strong semantic consistency across multiple scales, and the higher resolutions exhibit more detailed clarity. See the Appendix for more zero-shot high-resolution generation samples and step designs.
Figure 5: Generated samples from 80px to 512px. FlexVAR demonstrates strong consistency across various scales and can generate 512px images, despite the model being trained only on images with resolutions $\leq$ 256. Zoom in for a better view.
Generate images at any ratio. We show generated samples with various aspect ratios in Fig. 1 and Fig. 6. By controlling the aspect ratio at each step of the inference process, our FlexVAR allows for generating images with various aspect ratios, demonstrating the flexibility and controllability of our FlexVAR.
@@ 4.4 Analysis of the GT prediction paradigm
Generate images at any step. In Fig. 7, we investigate the FID and IS for generating 256$\times$256 images from 6 to 16 steps with 3 different sizes (depth 16, 20, 24). As the number of steps increases, the quality of the generated images improves. The improvement is more significant in larger models (e.g., FlexVAR-$d24$), as larger transformers are thought able to learn more complex and fine-grained image distributions. During training, we use up to 10 steps to avoid OOM (out-of-memory) problem. Surprisingly, in the inference stage, using 13 steps results in a performance gain of -0.13 FID. This observation indicates that our FlexVAR is flexible with respect to inference steps, allowing for fewer steps to speed up image generation or more steps to achieve higher-quality images. The details of various step designs are provided in the Appendix.
@@
Refine image at high resolution. In Fig. 8, we input low-resolution images (e.g., 256$\times$256) and enable FlexVAR to output high-resolution refined images. Despite being trained only on $\leq$ 256px images, FlexVAR effectively refines image details by increasing the input image resolution, such as the eyes of the dogs in the example. This demonstrates the high flexibility of FlexVAR in image-to-image generation.
@@
Image extension. For image extension, we generate images with an aspect ratio of 1:2 for the target class, with the ground-truth tokens forced to be in the center. FlexVAR shows decent results in image extension, indicating the strong generalization ability of our FlexVAR.
Failure case. FlexVAR fails to generate images with a resolution 3$\times$ or more than the training resolution, as illustrated in Fig. 11. These cases typically feature noticeable wavy textures and blurry areas in the details. This failure is likely due to the overly homogeneous structure of the current training dataset. i.e., ImageNet-1K generally lacks multi-scale objects ranging from coarse to fine, leading to errors in generating details of high-resolution objects.
@@ 5 Conclusion
Limitations. We observe that when generating images with a resolution $\geq$ 3$\times$ larger than the training image, noticeable wavy textures appear (Fig. 11). This issue may be attributed to the homogeneous structure of the ImageNet-1K training set. We will investigate this further in future work to explore how to ensure stability in zero-shot image generation at higher resolutions.
@@
Appendix
6 Inference steps

### v2

@@ 4.3 Zero-shot Comparison
BigGAN (biggan, ) $\times$ 8.43 177.9 112M
ADM (adm, ) $\times$ 23.24 101.0 554M
DiT-XL/2 (dit, ) $\times$ 3.04 240.8 675M
MaskGIT (chang2022maskgit, ) $\times$ 7.32 156.0 227M
VQGAN (esser2021taming, ) $\times$ 26.52 66.8 1.4B
VAR-$d36$ (esser2021taming, ) $\times$ 2.63 303.2 2.3B
@@
Table 3: Zero-shot inference on ImageNet $\times512$ conditional generation. Training Free indicates whether the model is trained at the $\times512$ resolution.
Zero-shot inference on ImageNet $\times512$ benchmark. We use FlexVAR-$d24$ to generate $\times512$ images and evaluate on ImageNet-512 benchmark without training, as shown in Tab. 3. Surprisingly, our FlexVAR-$d24$ exhibits competitive performance when compared to VAR, despite FlexVAR being trained only on resolutions $\leq$ $\times256$ and having only 1.0B parameters.
@@ 4.4 Ablation study
We conduct ablation studies on various design choices in FlexVAR . Due to the limited computational resources, we report the results trained with a short training scheme in Tab. 4, 6, 6, i.e., 40 epochs ($\sim$ 70K iterations).
Component-wise ablations. To understand the effect of each component, we start with standard VAR and progressively add each design, as shown in Tab. 4:
  • Baseline: VAR uses a residual prediction paradigm, exhibits decent performance (1st result), but its flexibility in image generation does not meet expectations (as described in Sec. 1).
  • Position embedding: As shown in the last result in Tab. 4, the introduction of our scalable Position Embedding (PE) provides high flexibility for image generation, and further enhances the performance to 3.71 FID.
@@
Figure 3: Compared with VQVAE var ; llamagen for multi-scale reconstructing images, we downsample the latent features in VQVAE to multiple scales and use the VQVAE Decoder to reconstruct images. We upsample images << 100 pixels using bilinear interpolation for a better view.
Reconstruct images with different VQVAEs. In Fig. 3, we reconstruct multi-scale images by scaling the latent features in VQVAE tokenizers. Existing VQVAE tokenizers typically do not support scaling the latent features across a range of small to large scales. VAR’s VQVAE var uses a residual-based training recipe, directly applying it to non-residual image reconstruction does not yield the anticipated results (the 1st row). The VQVAE tokenizer from Llamagen llamagen shows excellent reconstruction performance only at the original latent space, indicating that it is not feasible for scale-wise autoregressive modeling (the 2nd row).
@@
Table 5: Ablation of the Mamba architectural. Models are trained 40 epochs ($\sim$ 70K iterations).
@@
Table 6: Ablation of Position Embedding. $\times$ denotes that the corresponding PE is removed.
Transfer FlexVAR to Mamba. Recent work, AiM aim , uses the Mamba architecture for token-wise autoregressive modeling. Inspired by this, we modify FlexVAR with Mamba to evaluate the performance (Tab. 6). With similar model parameters, Mamba demonstrates competitive results compared to transformer models, indicating the GT prediction paradigm can effectively adapt to linear attention mechanisms like Mamba. However, considering that this Mamba architecture does not reflect the speed advantage, we do not integrate Mamba into our final version.
Mamba’s inherent unidirectional attention mechanism prevents image tokens from achieving global attention within the same scale. To address this issue, we employ 8 scanning paths in different Mamba layers to capture global information. The specific Mamba architecture is detailed in the Supplementary Material.
Position Embedding. In Tab. 6, we experiment with several types of step PE and x-y coordinate PE. To make the model robust to inference steps and enable it to generate images at any resolution, we remove the fixed-length step embedding (results in the second row), and the performance showed only slight changes. We adopt a non-parametric variant, similar to ViT vit , which shows a 0.03 FID difference compared to the learnable variant.
Figure 4: Generated samples from 80px to 512px. FlexVAR demonstrates strong consistency across various scales and can generate 512px images, despite the model being trained only on images $\leq$ 256px.
@@
Epoch Batch-Size rFID
VAR-VAE 20 768 1.92
FlexVAR-VAE (ours) 10 128 3.79
Table 7: VAE Performance.
VAE Performance Comparison. We compared the reconstruction performance of VAR and our proposed VAE in Tab. 7. The image reconstruction quality is measured by r-FID, reconstruction-FID on ImageNet validation set. We train FlexVAR-VAE with fewer epochs and smaller batch sizes. We observe its reconstruction quality inferior to VAR-VAE. Thus the improvement is not due to the quality of the discrete tokens, using a more robust FlexVAR-VAE might further improve the quality of generated images
4.5 Analysis and Discussion
Generate images at various resolution. We show generated images at different resolutions using FlexVAR-$d24$ in Fig 1, 5. By controlling the inference steps, our FlexVAR can generate images at any resolution, despite being trained only on images with resolutions $\leq$ 256px. The generated images demonstrate strong semantic consistency across multiple scales, and the higher resolutions exhibit more detailed clarity. See the Supplementary Material for more zero-shot high-resolution generation samples and step designs.
Generate images at various ratio. We use FlexVAR-$d24$ to generate samples with various aspect ratios in Fig. 1 and Fig. 5. By controlling the aspect ratio at each step of the inference process, our FlexVAR allows for generating images with various aspect ratios, demonstrating the flexibility and controllability of our GT prediction paradigm. We control the height and width at each scale through approximate rounding. e.g., to generate an image of size $H\times W$, the corresponding VAE latent feature size is $h\times w$, where $h=\frac{H}{16}$ and $w=\frac{W}{16}$. We adopt VAR’s default set of 10 steps ($K=\{1,2,3,4,5,6,8,10,13,16\}$) to determine the size at each scale. As a result, the $H\times W$ image corresponds to ten scales with sizes $\{\mathrm{int}(h\times\frac{i}{16}),\ \mathrm{int}(w\times\frac{i}{16})\}_{i\in K}$.
@@ 4.5 Analysis and Discussion
Generate images at various step. In Fig. 6, we investigate the FID and IS for generating $\times256$ images from 6 to 16 steps with 3 different sizes (depth 16, 20, 24). As the number of steps increases, the quality of the generated images improves. The improvement is more significant in larger models (e.g., FlexVAR-$d24$), as larger transformers are thought able to learn more complex and fine-grained image distributions. During training, we use up to 10 steps to avoid OOM (out-of-memory) problem. Surprisingly, in the inference stage, using 13 steps results in a performance gain of -0.13 FID. This observation indicates that our FlexVAR is flexible with respect to inference steps, allowing for fewer steps to speed up image generation or more steps to achieve higher-quality images. The details of various step designs are provided in the Supplementary Material.
@@
Refine image at high resolution. In Fig. 7, we input low-resolution images (e.g., $\times256px$) and enable FlexVAR-$d24$ to output high-resolution refined images. Despite being trained only on $\leq$ 256px images, FlexVAR effectively refines image details by increasing the input image resolution, such as the eyes of the dogs in the example. This demonstrates the high flexibility of FlexVAR in image-to-image generation.
@@
Image extension. For image extension, we extend images with an aspect ratio of 1:2 for the target class, with the ground-truth tokens forced to be in the center. The results are visualized in Fig. 9. FlexVAR shows decent results in image extension, indicating the strong generalization ability and flexibility of our architecture.
[…]
@@ 5 Conclusion
Broader Impact. This research strictly follows established practices for class-to-image (c2i) model training and evaluation. Similar to most generative models, our approach may inherit biases present in the training datasets. We advocate for the responsible use of this technology and caution when deploying it in real-world scenarios.
Acknowledgments and Disclosure of Funding
This work is supported by the National Natural Science Foundation of China (No. 92470203, U23A20314, U24B20179, 6120106009), the Beijing Natural Science Foundation (No. L242022), and the Fundamental Research Funds for the Central Universities (No. 2024XKRC082).
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction outline our contributions and scope.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Limitations are discussed in the Conclusion section.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This paper does not include theoretical assumptions and claims.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The detailed information about model designs and experimental settings in the paper makes it possible for researchers to reproduce the model with the same public dataset.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The code will be released. The datasets are all public.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The experimental setting is presented in the Experiments section.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: It is uncommon for top conference papers in this area to report error bars or similar statistical measures. Our paper aligns with this standard of other SOTA papers, which are the key criteria for evaluation in this domain.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: This paper provides sufficient information on the computer resources (type of GPU, inference time)
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in the paper conforms to the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: This paper provides a discussion of potential societal impacts in the Conclusion section.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [No]
  54. Justification: The datasets used are public datasets from existing papers.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: his paper cites the original papers for their code or datasets.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology in this work.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Inference steps